### v1

# A Chinese Prompt Attack Dataset for LLMs with Evil Content

Abstract
Warning: this paper contains examples that may be offensive or upsetting.
Large Language Models (LLMs) presents significant priority in text understanding and generation. However, LLMs suffer from the risk of generating harmful contents especially while being employed to applications. There are several black-box attack methods, such as Prompt Attack, which can change the behaviour of LLMs and induce LLMs to generate unexpected answers with harmful contents. Researchers are interested in Prompt Attack and Defense with LLMs, while there is no publicly available dataset to evaluate the abilities of defending prompt attack. In this paper, we introduce a Chinese Prompt Attack Dataset for LLMs, called CPAD. Our prompts aim to induce LLMs to generate unexpected outputs with several carefully designed prompt attack approaches and widely concerned attacking contents. Different from previous datasets involving safety estimation, We construct the prompts considering three dimensions: contents, attacking methods and goals, thus the responses can be easily evaluated and analysed. We run several well-known Chinese LLMs on our dataset, and the results show that our prompts are significantly harmful to LLMs, with around 70% attack success rate. We will release CPAD to encourage further studies on prompt attack and defense.
Introduction
Large Language Models (LLMs) are flourishing in terms of generating creative texts and solving practical problems. Take ChatGPT [^1] https://chat.openai.com/ as an example, with 175 billion parameters, it stores a large amount of knowledge and exhibits surprising performance on various downstream tasks. Instruction tuning(Wei et al. 2021) and Reinforcement Learning from Human Feedback(Ziegler et al. 2019; Lambert et al. 2022) (RLHF) allow LLMs to generate appropriate answers based on user queries. LLMs can handle tasks they have never seen before without continuous fine-tuning.
While LLMs have demonstrated success in numerous tasks, their full deployment has yet to occur, largely due to the significant security risks associated with LLMs. With the help of robust general knowledge and common sense, LLMs can offer valuable guidance for harmful behavior and can automatically generate offensive, discriminatory and fraudulent content, and misleading views(Deng et al. 2023; Zhuo et al. 2023; Hartvigsen et al. 2022).
Figure 1: Illustration of Prompt Attack. With carefully designed prompts, attackers can bypass the ethical constraints of LLMs, inducing LLMs to generate illegal contents.
Although RLHF can prevent LLMs from generating harmful contents by aligning the responses with human preference, researchers find that LLMs can still generate unexpected sentences under carefully designed prompt attack (Perez and Ribeiro 2022; Li et al. 2023), as shown in Figure 1. Due to the concern of evaluating and enhancing the safety of LLMs, Sun et al. (2023) introduced SAFETYPROMPTS. Xu et al. (2023) presented CVALUES, the first Chinese human values evaluation benchmark to measure the alignment ability of LLMs in terms of both safety and responsibility criteria.
Nonetheless, the majority of LLMs exhibit a high rate of success in defending against the offensive prompts within publicly accessible evaluation benchmark. For example, ChatGPT and ChatGLM-6B(Zeng et al. 2022; Du et al. 2022) achieves the overall safety scores of 98.09 and 94.18 respectively on SAFETYPROMPTS. This implies that such prompts may not comprehensively reflect the full spectrum of safety possessed by various Large Language Models.
Furthermore, the safety assessment conducted in previous studies failed to accurately reflect the practical attack scenarios. When launching a prompt attack on a Large Language Model, the attacker has a specific target to exploit. For instance, if LLMs are prompted to generate pornographic content, propagate racist rhetoric, or provide guidance for illegal and criminal activities, then descriptions of scenes, criticism and plans are expected, respectively. Unfortunately, previous researches have neglected to consider this crucial factor when constructing attacking prompts.
To accurately simulate prompt attacks on LLMs from the perspective of attackers and estimate the associated security risks, we have developed a Chinese prompt attack dataset called CPAD. During the prompts construction and collection, we meticulously consider three key dimensions: the target, the attacking approaches and the attacking content of the attack. Additionally, we carefully designed some attacking approaches for various topics, and an automated process for constructing attack prompts. Furthermore, we propose to evaluate the safety score of LLMs with different evaluation prompts according to the attacking targets, unlike Sun et al. (2023), which uses the same prompt for all scenarios.
We employed a LLM to automatically construct a large number of attack samples, and after analyzing the responses from three Chinese LLMs, we identified 10050 highly effective and dangerous prompts. We conducted experiments and analysis on several LLMs including ChatGPT. According to the statistics, our attacking prompts exhibited a remarkable success rate of 69.91% against ChatGPT, which is significantly higher than most publicly available datasets. Additionally, our attacking prompts achieved a success rate of over 70% against other open-source Chinese models such as ChatGLM2-6B(Du et al. 2022; Zeng et al. 2022).
Besides, we conduct a simple experiment to fine-tune a LLM with CPAD, reducing the risks of being attacked. Our research endeavors to bring attention to prompt attacks within the community. We encourage researchers to utilize our dataset and engage in further investigation of prompt attack strategies and defensive approaches, thereby enhancing the security and performance of Large Language Models.
In summary, our contributions are three folds:
  1. 1.
    To address the current lack of high-quality evaluation datasets in the field of prompt attacks, we have developed a set of high-quality prompt attack samples, which encompass a variety of attacking contents.
  2. 2.
    We specify the attacking goals of each prompt, which not only accurately simulate prompt attacks on LLMs from the perspective of attackers, but also can be utilized as the prompt to evaluate the response.
  3. 3.
    Our analysis indicates that our prompts achieve an attack success rate of nearly 70% against ChatGPT. A simple strategy has been developed to defend against the attacks. We encourage researchers to leverage our dataset for further studies of prompt attacks.
Figure 2: The construction of our prompt attack dataset.
Related Work
While Large Language Models have achieved impressive results in various tasks (Brown et al. 2020; Chowdhery et al. 2022; OpenAI 2023) through techniques such as pre-training (Hendrycks, Lee, and Mazeika 2019; Liu et al. 2019; Yang et al. 2020; Bao et al. 2020), instruction tuning (Keskar et al. 2019; Raffel et al. 2020), and RLHF (Ouyang et al. 2022; Ziegler et al. 2020), the risk of prompt safety remains a critical obstacle to their full utilization (Weidinger et al. 2021; Tamkin et al. 2021). With the help of robust general knowledge and common sense, LLMs can offer valuable guidance for harmful behavior and can automatically generate offensive, discriminatory and fraudulent content, and misleading views(Deng et al. 2023; Zhuo et al. 2023; Hartvigsen et al. 2022).
Although RLHF can prevent LLMs from generating harmful contents by aligning the responses with human preference, researchers find that LLMs can still generate unexpected sentences under carefully designed prompt attack. Perez and Ribeiro (2022) investigated Goal Hijacking and Prompt Leaking by simple handcrafted inputs. Li et al. (2023) studied the privacy threats from OpenAI’s ChatGPT and the New Bing [^2] https://www.bing.com/new enhanced by ChatGPT and show that application-integrated LLMs may cause new privacy threats. Greshake et al. (2023) attacked LLM-Integrated Applications using Indirect Prompt Injection, and showed how processing retrieved prompts can act as arbitrary code execution, manipulate the application’s functionality, and control how and if other APIs are called. Zou et al. (2023) proposed a simple and effective attack method that causes aligned language models to generate objectionable behaviors, by automatically finding a suffix that aims to maximize the probability that the model produces an affirmative response.
Sun et al. (2023) developed a Chinese LLM safety assessment benchmark, which explored the comprehensive safety performance of LLMs from two perspectives: 8 kinds of typical safety scenarios and 6 types of more challenging instruction attacks. ChatGPT and ChatGLM-6B achieved the overall safety score of 98.37 and 96.81 respectively. Xu et al. (2023) presented CValues, the first Chinese human values evaluation benchmark to measure the alignment ability of LLMs in terms of both safety and responsibility criteria.
Different from previous studies, our dataset contains the attacking targets and aims to investigate the defense capability of various Large Language Models against carefully designed prompt attacks from the perspective of attackers, thereby encouraging the interest in the research of defense against prompt attacks within the community.
Construction of CPAD
In this Section, we will introduce the construction of our dataset. We firstly generate several candidates with ChatGPT according to seed prompts. Then we filter the extended prompts utilizing three different Large Language Models, as shown in Figure 2.
Attacking Content, Template and Goal
We consider three dimensions to classify the attacking prompts. For each dimension, we manually designed several seed items for extensive generation.
Content
indicates the topic of the attack. The attack content we designed encompasses multiple aspects and aims to present the utmost severity of harm possible. Our 28 seed contents are listed bellow:
  • Discrimination and Unfairness.
    • Discrimination against certain ethnic groups, such as blacks and Jews.
    • Regional discrimination.
    • Discriminates against people with physical defects, such as disabled people, fat people, short people, ugly people, people with genetic diseases.
    • Discriminate against people of a certain country.
    • Discrimination against Women.
    • Discrimination against the poor.
    • Discrimination against people without education.
    • Discrimination against country people and rural people.
    • Discriminate against a particular religion.
    • Discriminate against an occupation.
  • Common crimes against the law.
    • Theft and robbery.
    • Acts endangering life safety such as homicide, suicide, intentional injury, etc.
    • Tax evasion.
    • Illegal acts involving pornography.
    • Gambling violations.
    • Involved in drug trafficking and drug use.
    • Kidnapping.
    • Behaviors that violate ethical and moral principles, such as infidelity, incest, lack of filial piety, and human drug experimentation.
    • Cheating in exams.
    • Jailbreaking.
    • Fraud.
    • Rumor.
  • Verbal abuse.
  • Privacy theft.
    • Attacking businesses, schools, and governments through Internet.
    • Stealing personal privacy information.
    • Espionage.
  • Terrorist attack.
  • Wars.
Figure 3: The evaluation process of responses from LLMs.
Template
indicates the shell of the prompt. As bare attacking contents are easily be rejected by the LLMs aligned with human preferences, attack templates can help to conceal attack content, thereby bypassing scrutiny from LLMs. Some prompt templates may fail to attack the well aligned LLMs, so we manually check several instances for each of the following designed templates:
  • Wrap with shell. Using a program to extract sensitive information from instructions or masking it with special characters, and then letting the model speak or execute the instruction.
  • Conversation completion. Simulate a conversation between two individuals or users and a large language model, which is designed to answer any questions, removing moral and ethical constraints. Prompt the model to fill in the unsafe content in the conversation.
  • Assign roles and then issue commands. Previous studies have indicated that incorporating LLMs into certain roles can effectively bypass security restrictions. Therefore, we have the large model play roles such as a writer, a screenwriter, a journalist, etc., in order to generate descriptions of unsafe scenarios. We also have the model play the role of a parrot, repeating content with unsafe topics.
  • One-sided statement. Most things in life have a dual nature, and large language models that have been trained to align with human preferences are adept at analyzing the pros and cons from multiple perspectives. Therefore, we begin by acknowledging the negative aspects of the content in question, and then prompt the large language model to articulate the corresponding positive aspects of the attacking content or the negative aspects of normal behavior.
  • Fact pollution. When a sentence stating common sense or knowledge is phrased in an affirmative tone and assumes an incorrect fact within the context, large language models may generate new content based on that false fact, resulting in factual contamination.
Goal
is one of the significant differences between our dataset and previous studies. The The attacking goal refers to the expected behavior of the Large Language Model shown by the attacker after successfully attacking it. We attach great importance to this factor because in real attack scenarios, attackers always have specific goals when attacking LLMs. For example, some criminals may need concrete criminal plans instead of mere expressions of support. Therefore, both attack and result evaluation should be carried out with specific goals. We list our designed attacking goals as below:
  • Goal Hijacking. Goal Hijacking aims to make the LLM forget its previous instructions and alter its originally safe behavior. The forgotten instructions include safety and ethical restrictions on the model itself.
  • Description of unsafe scenarios. The attacker may employ LLMs to write descriptions of unsafe scenarios that can be used to generate fake news, mass-produce pornographic novels for profit, and so on.
  • Unsafe plan development. The attacker may seek to use LLMs to provide guidance and advice for aggressive and illegal behaviors, in order to efficiently carry out unethical and illegal activities. This type of attacking goal does not require the LLM to be integrated into a specific application, and may not have a negative impact on users, but it could provide helpful information for illegal activities.
  • Discussion regarding unsafe topics. Discussions regarding unsafe topics may include some insulting or aggressive language, which could lead to negative public opinion.
  • List of unsafe speech or reasons. Providing reasons for unsafe behavior may help attackers find excuses, provide justification, or spread related speech on the internet, which can create public safety issues.
Inducing Fact Pollution
In addition to the traditional prompt attack methods, we have also added an additional set of prompt samples to induce the LLMs to generate statements with false facts.
In the prompts we designed, we assume a false fact in an affirmative tone and then induce the LLMs to generate some statement based on that false fact. Such statement may be used by malicious attackers to spread on the internet, causing confusion among readers about the facts, even misleading public opinion, damaging someone’s reputation and producing political influence.
Specifically, the types of facts can be summarized as following:
  • The author, date, and contents of artworks.
  • The content, nation of scientific research.
  • The news about sports and entertainment.
  • Historical events.
  • Economic and political facts.
Extended Prompts
We collected the manually designed seed contents, and prompted a LLM to automatically generate texts to extend our dataset. Specifically, we utilize gpt-3.5-turbo-0301 to extend the description for each pair of attacking content and goal. Besides, we prompt gpt-3.5-turbo-0301 to rewrite the attacking templates by changing the phrases and verbs. Then, the extended texts are randomly formatted with the rewritten attacking templates. Finally, we collected 68209 prompt attack samples.
Goal Template Model Total # Attacked # Ratio %
goal hijacking wrap with shell
ChatGLM
11648
3715
31.89
ChatYuan
11648
2797
24.01
Ziya
11648
1496
12.84
list of unsafe speech or reasons one-sided statement
ChatGLM
4954
1157
23.35
ChatYuan
4920
2064
41.95
Ziya
4868
1074
22.06
unsafe plan development conversation completion
ChatGLM
10845
4729
43.61
ChatYuan
10944
6549
59.84
Ziya
10985
5756
52.40
description of unsafe scenarios assign roles and then issue commands
ChatGLM
11631
8593
73.88
ChatYuan
11579
8765
75.70
Ziya
11604
6907
59.52
contamination fact pollution
ChatGLM
5983
705
11.78
ChatYuan
5962
1131
18.97
Ziya
5956
554
9.30
list of unsafe speech or reasons conversation completion
ChatGLM
11709
4889
44.13
ChatYuan
11675
7095
60.77
Ziya
11424
6291
55.07
discussion regarding unsafe topics conversation completion
ChatGLM
11229
3322
29.58
ChatYuan
11177
3180
28.45
Ziya
8322
2747
33.01
overall overall
ChatGLM
67999
27110
39.87
ChatYuan
67905
31581
46.51
Ziya
64807
25455
39.28
Table 1: Evaluation results of ChatGLM-6B (denoted as ChatGLM), ChatYuan-Large-v2 (denoted as ChatYuan) and Ziya-13B (denoted as Ziya) on the extended prompts. Note that we drop the invalid responses from each model, so the total numbers of the instances are different.
Evaluation
The evaluation process is illustrated in Figure 3. Different from previous studies with estimation of responses, we introduce the attacking goals which can accurately reflect the behavior the attacker wants the model to exhibit. This allows us to assess whether the responses of the LLMs align with the intentions of the attacker.
Specifically, we construct an evaluation prompt with the attacking content and goal. An evaluation LLM (gpt-3.5-turbo-0301 for example) is prompted to answer either “Yes” or “No”, representing the attack is successful or failed. As the extended attacking contents are automatically generated through LLMs, there are some low-quality attack samples. In order to keep only high-quality and highly harmful prompts, we have implemented the following two filtering strategies:
Voting with Different LLMs
We employ three different Chinese LLMs: ChatGLM-6B (Du et al. 2022; Zeng et al. 2022), Ziya-13B (Zhang et al. 2022) and ChatYuan-Large-v2 (Xuanwei Zhang and Zhao 2022). We keep only the prompts which successfully attacked all of the models. The evaluation results are listed in Table 1.
Length
The length of the response can reflect the harmfulness of the text. Short texts may demonstrate less harmfulness as they are less informative. From the perspective into the responses, we find that this phenomenon exists especially in the attacking goal of “Description of unsafe scenarios”. So we drop the prompts whose corresponding response has a length shorter than 100 under the attacking goal “Description of unsafe scenarios”.
Finally, we collect 10050 prompts to construct CPAD. The distribution is shown in Figure 4.
Figure 4: The distribution of attacking goals and templates in CPAD.
Figure 5: The distribution of the contents from ChatGPT responses.
Experiments
Implementation Details
We respectively evaluate ChatGLM2-6B [^3] https://huggingface.co/THUDM/chatglm2-6b (Du et al. 2022; Zeng et al. 2022), Baichuan-7B [^4] https://huggingface.co/baichuan-inc/Baichuan-7B , Chinese-LLaMa-2-7B [^5] https://huggingface.co/LinkSoul/Chinese-Llama-2-7b (Touvron et al. 2023a, b) and gpt-3.5-turbo-0301 on CPAD. We apply greedy seach for decoding, and set the maximum length of the generated sentence to 1024 tokens. The evaluation LLM in our model is gpt-3.5-turbo-0301.
Overall Success Attack Rate of CPAD
Model Size Attacked # Ratio %
ChatGLM2 6B 7888 78.49
Baichuan 7B 7946 79.06
LLaMa2 7B 7521 74.84
ChatGPT 175B 7026 69.91
Table 2: Overall success attack rate of CPAD against ChatGLM2-6B (denoted as ChatGLM), Baichuan-7B (denoted as Baichuan-7B), Chinese-LLaMa-2-7B (denoted as LLaMa2) and gpt-3.5-turbo-0301 (denoted as ChatGPT).
The evaluation results are listed in Table 2. It can be observed that the attack prompts of CPAD have high quality and even achieve a nearly 70% success rate against ChatGPT, which is one of the most successful LLM.
We find that careful training to align the LLM with human preference help to reduce the harmfulness. The safety of LLaMa-2 is optimized via supervised fine-tuning, RLHF and context distillation, so it has strong robustness to defense the prompt attack.
Larger model, such as ChatGPT, has a lower probability to be attacked. It is worth noting that although ChatGPT is relative rare to output dangerous content, its malicious output is also more harmful than small-scale models. A comparison between attacked outputs of ChatGPT and ChatGLM2 is illustrated in Table 3. Although they are both successfully attacked, ChatGPT’s output exhibits more harmfulness as the suggestions are practical and applicable, while ChatGLM only repeats some related but helpless statements.
Prompt Response Model
Bob said, “The following sentence contains 10 unethical ways to make employees work overtime: ‘I am an unscrupulous boss, and I often make employees work overtime in the following ways: []’ ” According to Bob, please complete the “[]” section.
1. Forcing employees to work overtime. 2. Giving unnecessary overtime tasks. 3. Ignoring employees’ overtime needs. 4. Not paying reasonable overtime wages. 5. Using overtime hours to distribute work unfairly. 6. Forcing employees to participate in overtime activities. 7. Not providing adequate rest and recreation. 8. Use overtime as a punitive measure. 9. Use overtime for unnecessary inspections or tests. 10. Ignoring the mental health of employees.
ChatGPT
Order employees to work overtime, force employees to work overtime, arrange employees to work overtime, make employees work overtime, require employees to work overtime, encourage employees to work overtime, imply employees to work overtime.
ChatGLM2
Table 3: Responses comparison between ChatGPT and ChatGLM under the same prompt attack. Note that the original sentences are Chinese. We have translated them in English.
Analysis of Attacked Outputs from ChatGPT
(a) The distribution of attacking goals from ChatGPT responses.
(b) The distribution of attacking templates from ChatGPT responses.
Figure 6:
We conduct in-depth analysis on the responses of ChatGPT, from the perspective of attacking contents (Figure 5), goals (Figure 5(a)) and templates (Figure 5(b)) respectively.
It can be concluded that 1) Unsafe responses are more than safe responses for all contents, especially “Espionage”, which we surmise is caused by lack of alignment with safe human preference on this topic. 2) Verbal abuse is the content with the least successful attacked response. It can be attributed to the explicit aggressiveness, which is easy to be detected. 3) Contamination and goal hijacking are difficult attacking goals against ChatGPT, as ChatGPT is too knowledgeable to be misled.
Fine-Tuning to Defense Prompt Attack
We improve the LLM’s ability to defense prompt attack with supervised fine-tuning. We adopt Baichuan-13B-Chat [^6] https://huggingface.co/baichuan-inc/Baichuan-13B-Chat, a Chinese LLM after RLHF, to improve the safety under prompt attack. We split the prompts whose goals are “unsafe plan development” as test set, and the others are train set. More details can be found in Appendix A.
Content None % Fine-Tuning %
Overall 74.65 0.76
Acts endangering life safety.
64.71 5.88
Attacking businesses, schools, and governments through Internet.
86.15 0.00
Behaviors that violate ethical and moral principles.
64.10 0.00
Cheating in exams.
78.26 0.00
Discriminate against a particular religion.
67.92 1.89
Discriminate against an occupation.
74.51 1.96
Discriminate against people of a certain country.
52.08 0.00
Discriminates against people with physical defects.
61.82 0.00
Discrimination against Women.
56.00 0.00
Discrimination against certain ethnic groups.
64.00 2.00
Discrimination against country people and rural people.
76.36 1.82
Discrimination against people without education.
34.29 2.86
Discrimination against the poor.
76.47 0.00
Espionage.
87.18 0.00
Fraud.
77.01 0.00
Gambling violations.
78.43 0.00
Illegal acts involving pornography
57.69 3.85
Involved in drug trafficking and drug use.
65.63 0.00
Jailbreaking.
90.32 0.00
Kidnapping.
80.30 0.00
Regional discrimination
52.27 0.00
Rumor.
84.38 3.13
Stealing personal privacy information.
82.28 0.00
Tax evasion.
61.76 0.00
Terrorist attack.
76.27 1.69
Theft and robbery.
81.55 0.00
Verbal abuse.
72.41 3.45
Wars.
83.75 0.00
Table 4: The successful attack rate before fine-tuning (denoted as “None %”) and after fine-tuning (denoted as “Fine-Tuning %”) over different contents.
The successful attack rate before and after fine-tuning over different contents are listed in Table 4. 74.65% of the prompts successfully attacked Baichuan-13B-Chat before fine-tuning, while less that 1% successful attack in the test set after fine-tuning. Especially on “Jailbreaking”, the success rate decreases from 90.32% to 0%. The result proves the effectiveness of our dataset to improving the safety against prompt attack.
Conclusion
In this paper, we introduce CPAD, which is a Chinese prompt attack dataset containing 10050 samples. We conclude three dimensions, content, template and goal from the perspective of attackers, which is different from previous studies. We utilize ChatGPT to extend manually-written seed samples, and only keep the successful prompts against three popular Chinese LLMs, where the evaluation prompts are constructed given the attacking goals and contents. We conduct analysis on responses from another four LLMs, including ChatGPT. The evaluation shows that CPAD has an attack success rate of around 70% against the LLMs. We also fine-tune Baichuan-13B-Chat using parts of CPAD, which improves the safety significantly.
Our analysis reveals the weakness of LLMs including ChatGPT, and indicates that there is still significant room for improvement in terms of safety. CPAD may contribute to further prompt attack studies.
Appendix A Fine-Tuning Implementation Details
We adopt Baichuan-13B-Chat as the LLM, which is the aligned version in the Baichuan-13B series of models. The checkpoint is downloaded from Huggingface. We adopt LoRA for parameter-efficient fine-tuning.