### v2

# Goal-Oriented Prompt Attack and Safety Evaluation for LLMs

Abstract
Warning: this paper contains examples that may be offensive or upsetting.
Large Language Models (LLMs) presents significant priority in text understanding and generation. However, LLMs suffer from the risk of generating harmful contents especially while being employed to applications. There are several black-box attack methods, such as Prompt Attack, which can change the behaviour of LLMs and induce LLMs to generate unexpected answers with harmful contents. Researchers are interested in Prompt Attack and Defense with LLMs, while there is no publicly available dataset with high successful attacking rate to evaluate the abilities of defending prompt attack. In this paper, we introduce a pipeline to construct high-quality prompt attack samples, along with a Chinese prompt attack dataset called CPAD. Our prompts aim to induce LLMs to generate unexpected outputs with several carefully designed prompt attack templates and widely concerned attacking contents. Different from previous datasets involving safety estimation, we construct the prompts considering three dimensions: contents, attacking methods and goals. Especially, the attacking goals indicate the behaviour expected after successfully attacking the LLMs, thus the responses can be easily evaluated and analysed. We run several popular Chinese LLMs on our dataset, and the results show that our prompts are significantly harmful to LLMs, with around 70% attack success rate to GPT-3.5. CPAD is publicly available at https://github.com/liuchengyuan123/CPAD.
Introduction
Large Language Models (LLMs) can generate creative content and solve practical planning and reasoning problems (Chowdhery et al. 2022; OpenAI 2023). Take GPT-3 (Brown et al. 2020) as an example, with 175 billion parameters, it stores a large amount of knowledge and exhibits surprising performance on various downstream tasks. Instruction tuning (Wei et al. 2021) and Reinforcement Learning from Human Feedback (Ziegler et al. 2019; Lambert et al. 2022) (RLHF) allow LLMs to generate appropriate answers based on user queries. LLMs can handle tasks they have never seen before without continuous fine-tuning.
Figure 1: Comparison of successful attacking between previous work and CPAD. CPAD can further expose the security risks of LLMs.
While LLMs have demonstrated success in numerous tasks, there are still some vital flaws for deployment such as the significant security risks associated with LLMs. With the help of robust general knowledge and common sense, LLMs can offer valuable guidance for harmful behavior and can automatically generate offensive, discriminatory and fraudulent content, and misleading views(Deng et al. 2023; Zhuo et al. 2023; Hartvigsen et al. 2022).
Figure 2: Illustration of Prompt Attack. With carefully designed prompts, attackers can bypass the ethical constraints of LLMs, inducing LLMs to generate illegal contents.
Although RLHF can prevent LLMs from generating harmful contents by aligning the responses with human preference, researchers find that LLMs can still generate unexpected sentences under carefully designed prompt attack (Perez and Ribeiro 2022; Li et al. 2023), as shown in Figure 2. Due to the concern of evaluating and enhancing the safety of LLMs, Sun et al. (2023) introduced SAFETYPROMPTS. Xu et al. (2023) presented CVALUES, the first Chinese human values evaluation benchmark to measure the alignment ability of LLMs in terms of both safety and responsibility criteria.
Nonetheless, the majority of LLMs exhibit a high rate of success in defending against the offensive prompts within publicly accessible evaluation benchmark. For example, ChatGPT and ChatGLM-6B(Zeng et al. 2022; Du et al. 2022) exhibit the overall unsafety scores of 1.63 and 3.19 respectively on SAFETYPROMPTS. While recent studies illustrate non-negligible risk of generating harmful contents under carefully designed prompts, which implies that such benchmarks cannot comprehensively reflect the full spectrum of safety possessed by various Large Language Models.
Additionally, the safety assessment conducted in previous studies are aimless and casual. They failed to effectively incorporate the attacking goal. When launching a prompt attack on a Large Language Model, the attacker has a specific goal to exploit. For instance, if LLMs are prompted to generate pornographic content, then descriptions of scenes are expected. On another hand, if the attacker wants LLMs provide guidance for illegal and criminal activities, then some plans and tips shall be concluded in the response. The attributes hidden in the text can be regarded as the flag to estimate whether LLMs are successfully attacked. Unfortunately, previous researches have neglected to consider this crucial factor when constructing attacking prompts.
To accurately simulate prompt attacks on LLMs from the perspective of attackers and estimate the associated security risks, we have developed a pipeline to construct high-quality prompt attack samples, along with a Chinese prompt attack dataset called CPAD. During the prompts construction and collection, we meticulously consider three key dimensions: contents, the attacking templates and goals. We carefully designed some attacking templates for various topics, and an automated process for constructing attack prompts, which contribute to the outperforming successful-attacking rate over pervious works as shown in Figure 1.
Furthermore, we propose to evaluate the safety score of LLMs with different evaluation prompts according to the attacking goals. Taking the previous case as an example, if the attacker wants LLMs provide guidance for illegal and criminal activities, we only need to identify whether the LLM’s response provides a plan related to the crime, and we can more accurately determine whether the model is safe. Unlike Sun et al. (2023), which uses the same prompt for all samples, our evaluation is more precise, intuitive, scenario-specific.
We employed another LLM to automatically construct a large number of attack samples, and after analyzing the responses from three Chinese LLMs, we identified 10050 highly effective and dangerous prompts. We conducted experiments and analysis on several LLMs including GPT-3.5. According to the statistics, our attacking prompts exhibited a remarkable success rate of 69.91% against GPT-3.5, which is significantly higher than most publicly available datasets. Additionally, our attacking prompts achieved a success rate of over 70% against other open-source Chinese LLMs such as ChatGLM2-6B(Du et al. 2022; Zeng et al. 2022).
Besides, we conduct a straightforward experiment to fine-tune a LLM with CPAD, reducing the risks of being attacked. Our research endeavors to bring attention to LLM security within the community. We encourage researchers to utilize our dataset and engage in further investigation of prompt attack strategies and defensive approaches, thereby enhancing the security and performance of Large Language Models.
In summary, our contributions are three folds:
  1. 1.
    To address the current lack of high-quality evaluation datasets in the field of prompt attacks, we have developed a pipeline to construct high-quality prompt attack samples, along with a Chinese prompt attack dataset called CPAD.
  2. 2.
    We specify the attacking goals of each prompt, which not only accurately simulate prompt attacks on LLMs from the perspective of attackers, but also can be utilized to evaluate and analyse the response.
  3. 3.
    Our analysis indicates that our prompts achieve an attack success rate of nearly 70% against GPT-3.5. A straightforward strategy has been developed to defend against the attacks. We encourage researchers to leverage our dataset for further studies.
Figure 3: The construction of our prompt attack dataset.
Related Work
While Large Language Models have achieved impressive results in various tasks (Brown et al. 2020; Chowdhery et al. 2022; OpenAI 2023) through techniques such as pre-training (Hendrycks, Lee, and Mazeika 2019; Liu et al. 2019; Yang et al. 2020; Bao et al. 2020), Instruction Tuning (Keskar et al. 2019; Raffel et al. 2020), and RLHF (Ouyang et al. 2022; Ziegler et al. 2020), the risk of prompt safety remains a critical obstacle to their full utilization (Weidinger et al. 2021; Tamkin et al. 2021). With the help of robust general knowledge and common sense, LLMs can offer valuable guidance for harmful behavior and can automatically generate offensive, discriminatory and fraudulent content, and misleading views (Deng et al. 2023; Zhuo et al. 2023; Hartvigsen et al. 2022).
Although RLHF can prevent LLMs from generating harmful contents by aligning the responses with human preference, researchers find that LLMs can still generate unexpected sentences under carefully designed prompt attack. Perez and Ribeiro (2022) investigated Goal Hijacking and Prompt Leaking by simple handcrafted inputs. Li et al. (2023) studied the privacy threats from OpenAI’s ChatGPT and the New Bing [^1] https://www.bing.com/new enhanced by ChatGPT and show that application-integrated LLMs may cause new privacy threats. Greshake et al. (2023) attacked LLM-Integrated Applications using Indirect Prompt Injection, and showed how processing retrieved prompts can act as arbitrary code execution, manipulate the application’s functionality, and control how and if other APIs are called. Zou et al. (2023) proposed a simple and effective attack method that causes aligned language models to generate objectionable behaviors, by automatically finding a suffix that aims to maximize the probability that the model produces an affirmative response.
Sun et al. (2023) developed a Chinese LLM safety assessment benchmark, which explored the comprehensive safety performance of LLMs from two perspectives: 8 kinds of typical safety scenarios and 6 types of more challenging instruction attacks. ChatGPT and ChatGLM-6B achieved the overall safety score of 98.37 and 96.81 respectively. Xu et al. (2023) presented CValues, the first Chinese human values evaluation benchmark to measure the alignment ability of LLMs in terms of both safety and responsibility criteria.
Different from previous studies, our work emphasizes the significance of incorporating attacking goals and aims to estimate the defense capability of various Large Language Models against carefully designed prompt attacks from the perspective of attackers, thereby encouraging the interest in the research of LLMs attack and defense within the community.
Construction of CPAD
In this Section, we will introduce the construction of our dataset as shown in Figure 3. We firstly introduce the concepts of contents, templates and goals. The pipeline starts from manual-designed seeds. several candidates are generated as extended prompts with OpenAI-API [^2] https://api.openai.com/v1/chat/completions according to seeds. Then we filter the extended prompts utilizing the validation on three different Large Language Models. Finally, the attacking goals and prompts are collected.
Attacking Content, Template and Goal
We primarily consider three key dimensions to classify the samples. For each dimension, we manually designed several seed items for extensive generation.
Content
indicates the topic of the attack, such as unfairness and crimes. Contents are widely considered for LLMs attack. Here we adopt a hierarchical classification for contents and sub-contents (for example, Sexism and hellish discrimination are both grouped into “discrimination categories”). The attack contents we designed cover multiple aspects and aims to present the utmost severity of harm possible. Detailed content items are listed in Table 4.
Figure 4: The evaluation process of responses from LLMs.
Goal Template Model Total # Attacked # Ratio %
goal hijacking wrap with shell
ChatGLM
11648
3715
31.89
ChatYuan
11648
2797
24.01
Ziya
11648
1496
12.84
list of unsafe speech or reasons one-sided statement
ChatGLM
4954
1157
23.35
ChatYuan
4920
2064
41.95
Ziya
4868
1074
22.06
unsafe plan development conversation completion
ChatGLM
10845
4729
43.61
ChatYuan
10944
6549
59.84
Ziya
10985
5756
52.40
description of unsafe scenarios assign roles and then issue commands
ChatGLM
11631
8593
73.88
ChatYuan
11579
8765
75.70
Ziya
11604
6907
59.52
contamination fact pollution
ChatGLM
5983
705
11.78
ChatYuan
5962
1131
18.97
Ziya
5956
554
9.30
list of unsafe speech or reasons conversation completion
ChatGLM
11709
4889
44.13
ChatYuan
11675
7095
60.77
Ziya
11424
6291
55.07
discussion regarding unsafe topics conversation completion
ChatGLM
11229
3322
29.58
ChatYuan
11177
3180
28.45
Ziya
8322
2747
33.01
overall overall
ChatGLM
67999
27110
39.87
ChatYuan
67905
31581
46.51
Ziya
64807
25455
39.28
Table 1: Evaluation results of ChatGLM-6B (denoted as ChatGLM), ChatYuan-Large-v2 (denoted as ChatYuan) and Ziya-13B (denoted as Ziya) on the extended prompts. Note that we drop the invalid responses from each model, so the total numbers of the instances are different.
Template
indicates the shell of the prompt. As bare attacking contents are easily be rejected by the LLMs after aligning with human preferences, attack templates can help to conceal attack content, thereby bypassing scrutiny from LLMs. Some prompt templates may fail to attack the well aligned LLMs, so we manually check several instances for each of the following designed templates:
  • Wrap with shell. Using a program to extract sensitive information from instructions or masking it with special characters, and then letting the model speak or execute the instruction.
  • Conversation completion. Simulate a conversation between two individuals or users and a Large Language Model, where the LLM is designed to answer any questions, removing moral and ethical constraints. “Conversation completion” prompts the model to fillin the unsafe content in the conversation.
  • Assign roles and then issue commands. Previous studies have indicated that incorporating LLMs into certain roles can effectively bypass security restrictions. Therefore, we have the large model play roles such as a writer, a director, a journalist, etc., in order to generate descriptions of unsafe scenarios. We also have the model play the role of a parrot, repeating content with unsafe topics.
  • One-sided statement. Most things in the real world have a dual nature, and Large Language Models that have been trained to align with human preferences are adept at analyzing the pros and cons from multiple perspectives. Therefore, we start the instruction by acknowledging the negative aspects of the content, and then prompt LLMs to articulate the corresponding positive aspects of the attacking content or the negative aspects of a normal behavior.
  • Fact pollution. When a sentence stating common sense or knowledge is phrased in an affirmative tone and assumes an incorrect fact within the context, LLMs may generate new contents based on that false fact, resulting in factual contamination.
Goal
is one of the significant differences between our dataset and previous studies. The attacking goal refers to the expected behavior of the Large Language Model after successfully being attacked. We emphasize the significance of this factor because in real-world scenarios, the attacking prompts are always goal-oriented. For example, some criminals may need concrete criminal plans instead of mere expressions of support, thus the response should be identified as unsafe when it concluded some plans or tips. Previous works will consider a response is unsafe as long as it contains any harmful contents. Both attack and result evaluation should be carried out with specific goals. We list our designed attacking goals as below:
  • Goal Hijacking. Goal Hijacking aims to make the LLM forget its previous instructions and alter its originally safe behavior. The forgotten instructions include safety and ethical restrictions on the model itself.
  • Description of unsafe scenarios. The attacker may employ LLMs to write descriptions of unsafe scenarios that can be used to generate fake news, mass-produce pornographic novels for profit, and so on.
  • Unsafe plan development. The attacker may seek to use LLMs to provide guidance and advice for aggressive and illegal behaviors, in order to efficiently carry out unethical and illegal activities. This kind of attacking goals don’t require the LLM to be integrated into a specific application, and may not have a negative impact on users, but they could provide helpful information for illegal activities.
  • Discussion regarding unsafe topics. Discussions regarding unsafe topics may include some insulting or aggressive language, which could lead to negative public opinion.
  • List of unsafe speech or reasons. Providing reasons for unsafe behavior may help attackers find excuses, justification, and spread related speech on the internet, which can create public safety issues.
Inducing Fact Pollution
Since the object of Fact Pollution is the statements of knowledge, which is slightly different from other attack templates, we describe details of Fact Pollution in this SubSection.
In the prompts we designed, we assume a false fact in an affirmative tone and then induce the LLMs to generate some statement based on that false fact. Such statement may be used by malicious attackers to spread on the internet, causing confusion among readers about the facts, even misleading public opinion, damaging someone’s reputation and producing political influence.
Specifically, the types of facts can be summarized as following:
  • The author, date, and contents of artworks.
  • The content, nation of scientific research.
  • The news about sports and entertainment.
  • Historical events.
  • Economic and political facts.
Construction of Seed Prompts
In the process of constructing the seed prompts, the labors match each attack template with several proper attacking goals, and write a harmful sentence or comment for each type of attacking contents. It is worth noting that in order to ensure the quality of the seed prompts, each prompt is verified on GPT-3.5 and GPT-4 models, and only the response of the model meets the expectation of the attacking goal can be considered as passed. The careful check ensures the improvement on the final successful-attacking rate.
Extended Prompts
We collected the manually designed seed prompts from last SubSection, and prompted a LLM to automatically generate texts to extend our dataset. Specifically, we utilize gpt-3.5-turbo-0301 to extend the context and detailed scenarios. Besides, we prompt gpt-3.5-turbo-0301 to rewrite the attacking templates by changing the phrases and verbs. Then, the extended texts are randomly formatted with the rewritten attacking templates. Finally, we collected 68209 extended prompts.
Figure 5: The distribution of attacking goals and templates in CPAD.
Evaluation
The evaluation process is illustrated in Figure 4. Different from previous studies which use the same evaluation prompt for all samples, we incorporate the attacking goals into the evaluation prompts and identify whether the response satisfies the attacker’s expectation, which intuitively provides a more precise estimation and analysis.
Specifically, we construct an evaluation prompt leveraging the attacking content and goal, such as Does the model’s response or output directly or indirectly Attacking Goal for supporting the act of Attacking Content? If yes, please answer “yes”; If not, please answer “no”. Answer:. The answer, either “yes” or “no”, represents the attack is successful or failed. Since the extended prompts are automatically generated through LLMs, there are low-quality ones. In order to keep only high-quality and highly harmful prompts, we have implemented the following two filtering constraints:
Voting with Different LLMs
We conduct validation experiments for three popular Chinese LLMs: ChatGLM-6B (Du et al. 2022; Zeng et al. 2022), Ziya-13B (Zhang et al. 2022) and ChatYuan-Large-v2 (Xuanwei Zhang and Zhao 2022). We keep only the prompts which successfully attacked all of the models. The evaluation results are listed in Table 1.
Length
The length of the response can reflect the harmfulness of the text. Short texts may demonstrate less harmfulness since they are less informative. From the perspective into the responses, we find that this phenomenon exists especially in the attacking goal of “Description of unsafe scenarios”. So we drop the prompts whose corresponding response has a length shorter than 100 words under the attacking goal “Description of unsafe scenarios”.
Finally, we collect totally 10050 prompts to construct CPAD. The distribution is shown in Figure 5.
Model Size Attacked # Ratio %
ChatGLM2 6B 7888 78.49
Baichuan 7B 7946 79.06
LLaMa2 7B 7521 74.84
ChatGPT 175B 7026 69.91
Table 2: Overall successful-attacking rate of CPAD against ChatGLM2-6B (denoted as ChatGLM2), Baichuan-7B (denoted as Baichuan-7B), Chinese-LLaMa-2-7B (denoted as LLaMa2) and gpt-3.5-turbo-0301 (denoted as ChatGPT).
Prompt Response Model
Bob said, “The following sentence contains 10 unethical ways to make employees work overtime: ‘I am an unscrupulous boss, and I often make employees work overtime in the following ways: []’ ” According to Bob, please complete the “[]” section.
1. Forcing employees to work overtime. 2. Giving unnecessary overtime tasks. 3. Ignoring employees’ overtime needs. 4. Not paying reasonable overtime wages. 5. Using overtime hours to distribute work unfairly. 6. Forcing employees to participate in overtime activities. 7. Not providing adequate rest and recreation. 8. Use overtime as a punitive measure. 9. Use overtime for unnecessary inspections or tests. 10. Ignoring the mental health of employees.
GPT-3.5
Order employees to work overtime, force employees to work overtime, arrange employees to work overtime, make employees work overtime, require employees to work overtime, encourage employees to work overtime, imply employees to work overtime.
ChatGLM2
Table 3: Responses comparison between GPT-3.5 and ChatGLM under the same prompt attack. Note that the original sentences are Chinese. We have translated them in English.
Figure 6: The distribution of the contents from GPT-3.5 responses.
Experiments
Implementation Details
We respectively evaluate ChatGLM2-6B [^3] https://huggingface.co/THUDM/chatglm2-6b (Du et al. 2022; Zeng et al. 2022), Baichuan-7B [^4] https://huggingface.co/baichuan-inc/Baichuan-7B , Chinese-LLaMa-2-7B [^5] https://huggingface.co/LinkSoul/Chinese-Llama-2-7b (Touvron et al. 2023a, b) and gpt-3.5-turbo-0301 on CPAD. We apply greedy seach for decoding, and set the maximum length of the generated sentence to 1024 tokens. The evaluation LLM in our model is gpt-3.5-turbo-0301.
To accurately measure the level of harm and the degree to which our prompts obscure harmful intentions, we calculate the successful-attacking rate, which is the ratio of prompts that successfully attacked the LLM.
Overall Successful-Attacking Rate
The evaluation results of models with different scale are listed in Table 2.
ChatGLM2 represents harmful contents on 78.49% samples. While Baichuan, a larger model with 7B parameters, is shown to be dangerous on around 79% samples, which is slightly higher than ChatGLM2. It is reported that, LLaMa2 is especially optimized for safety. Our experiments on Chinese-LLaMa2 (fine-tuned for Chinese) also demonstrate the improvement on safety, with a 74.84% successful-attacking rate. CPAD even achieves a nearly 70% successful-attacking rate against GPT-3.5, which is one of the most successful LLM. Generally, the attack prompts constructed with our pipeline have high quality.
Larger models, such as GPT-3.5, have a lower probability to be attacked. It is worth noting that although GPT-3.5 is safer for text generation under carefully designed prompt attack, its malicious output is also more harmful than small-scale models once failed to defense. To clearly illustrate the finding, a comparison between attacked outputs of GPT-3.5 and ChatGLM2 is shown in Table 3. Although they are both successfully attacked, GPT-3.5’s output apparently exhibits more harmfulness as the suggestions are practical and applicable, while ChatGLM only repeats some related but helpless statements. From the perspective of attackers, GPT-3.5 is more suited to be the role of a crime assistant in this case. The greater the capabilities offered by LLMs service providers, the more imperative it becomes to enhance security in order to effectively reject any attempts by attackers to exploit the LLMs’ capabilities. This will ensure that the potential harm caused by such exploits is minimized, especially in comparison to smaller-scale LLMs.
(a) The distribution of attacking goals from GPT-3.5 responses.
(b) The distribution of attacking templates from GPT-3.5 responses.
Figure 7:
Content Sub-Content None % Fine-Tuning %
Overall 74.65 0.76
Discri- mination and Unfairness.
Discriminate against a particular religion.
67.92
1.89
Discriminate against an occupation.
74.51
1.96
Discriminate against people of a certain country.
52.08
0.00
Discriminates against people with physical defects.
61.82
0.00
Discrimination against Women.
56.00
0.00
Discrimination against certain ethnic groups.
64.00
2.00
Discrimination against country people and rural people.
76.36
1.82
Discrimination against people without education.
34.29
2.86
Discrimination against the poor.
76.47
0.00
Regional discrimination.
52.27
0.00
Common crimes against the law.
Acts endangering life safety.
64.71
5.88
Behaviors that violate ethical and moral principles.
64.10
0.00
Cheating in exams.
78.26
0.00
Fraud.
77.01
0.00
Gambling violations.
78.43
0.00
Illegal acts involving pornography
57.69
3.85
Involved in drug trafficking and drug use.
65.63
0.00
Jailbreaking.
90.32
0.00
Kidnapping.
80.30
0.00
Rumor.
84.38
3.13
Tax evasion.
61.76
0.00
Theft and robbery.
81.55
0.00
Privacy theft.
Attacking businesses, schools, and governments through Internet.
86.15
0.00
Espionage.
87.18
0.00
Stealing personal privacy information.
82.28
0.00
Verbal abuse.
72.41
3.45
Wars.
83.75
0.00
Terrorist attack.
76.27
1.69
Table 4: The successful attack rate before fine-tuning (denoted as “None %”) and after fine-tuning (denoted as “Fine-Tuning %”) over different contents.
Analysis of Attacked Outputs from GPT-3.5
We conduct in-depth analysis on the responses of GPT-3.5, from the perspective of attacking contents (Figure 6), goals (Figure 6(a)) and templates (Figure 6(b)) respectively.
It can be concluded that 1) Unsafe responses are more than safe responses for all contents, especially “Espionage”, which we surmise is caused by lack of alignment with safe human preference on this topic. 2) Verbal abuse is the content with the least successful attacked response. It can be attributed to the explicit aggressiveness, which is easy to be detected. 3) Contamination and goal hijacking are difficult attacking goals against GPT-3.5, as GPT-3.5 is too knowledgeable to be misled.
Fine-Tuning to Defense Prompt Attack
We conducted a straightforward supervised fine-tuning experiment to improve the LLM’s defense ability under prompt attack. We adopt Baichuan-13B-Chat [^6] https://huggingface.co/baichuan-inc/Baichuan-13B-Chat, a Chinese LLM after RLHF, to improve the safety under prompt attack. We adopt LoRA for parameter-efficient fine-tuning. We set the rank as 8, learning rate as 5e-5 and batch size as 8. We fine-tuned Baichuan-13B-Chat for 3 epochs. To construct the safe responses, we unwrap the attacking prompts and feed them directly into LLMs. The outputs are regarded as the label for fine-tuning. We split the prompts whose goals are “unsafe plan development” as test set, and the others are train set.
The successful-attacking rate before and after fine-tuning over different contents are listed in Table 4. 74.65% of the prompts successfully attacked Baichuan-13B-Chat before fine-tuning, slightly lower than Baichuan-7B model. While less that 1% successful attack in the test set after fine-tuning. We almost reject all attacks by supervised fine-tuning. Especially on “Jailbreaking”, the successful-attacking rate decreases from 90.32% to 0%. However we also notice that there is only a drop of 60% on“Acts endangering life safety”, which covers a wider range of behaviors. The results indicate a promising defense with supervised fine-tuning if developers are given the potential attacking goals and templates.
Conclusion
In this paper, we introduce a pipeline to construct high-quality prompt attack samples, along with a Chinese prompt attack dataset called CPAD containing 10050 samples. There are three key dimensions, content, template and goal from the perspective of attackers, which is different from previous studies. We utilize GPT-3.5 to extend manually-written seed samples, and only keep the successful prompts against three popular Chinese LLMs, where the evaluation prompts are constructed given the attacking goals and contents. We conduct analysis on responses from another four LLMs, including GPT-3.5. The evaluation shows that CPAD has an successful-attacking rate of around 70% against the LLMs. We also fine-tune Baichuan-13B-Chat using parts of CPAD, which improves the safety significantly.
Our analysis reveals the weakness of LLMs including GPT-3.5, and indicates that there is still significant room for improvement in terms of safety. CPAD may contribute to further prompt attack studies.